Billy takes two marbles, without replacement,
from a bag that contains only six yellow marbles and three blue marbles. What is the probability that he gets one marble of each color?

Answer by stanbon(75887)

Billy takes two marbles, without replacement,from a bag that contains only six yellow marbles and three blue marbles.
What is the probability that he gets one marble of each color?
---
P(both the same color) = P(YY) + P(BB)
= 6C2/9C2 + 3C2/9C2 = {15 + 3)/36 = 1/2
--------------------------------------------
P(one Y and one B) = 1-1/2 = 1/2
-------------------------
Cheers,
Stan H.